OpenStudy (anonymous):

evaluate the value of A = ? A = 1^3 - 2^3 + 3^3 - 4^3 + 5^3 - 6^3 + ... + 99^3

OpenStudy (anonymous):

Okay well there is an easy way to do this problem and there is also a hard way. Let me look in my Calculus notes to help you out with this problem.

OpenStudy (anonymous):

Okay here we go. Your problem comes down to Sigma E = i^3 where i =1 and n = 99. You can figure this out with a formula, which i will give you in a second.

OpenStudy (anonymous):



OpenStudy (anonymous):

still confused what to do next ?

OpenStudy (anonymous):

It's n^2(n+1)^2/4

OpenStudy (anonymous):

This is your formula you need to use, where n = 99

OpenStudy (anonymous):

99^2(100)^2/4 = your answer

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

how do i know it would be equals n^2(n+1)^2/4 ?

OpenStudy (anonymous):

This solution is cheating in a way. It is like magic... Formulas appearing from thin air

OpenStudy (anonymous):

Because the exponent is 3. If you look at the picture I gave the link to, the other formulas are for if the exponent is 2, 1, or even a constant.

OpenStudy (anonymous):

The formula is defined in Summation rules for Sigma Notation

OpenStudy (anonymous):

the 4th formula is when all series positive, while the question above is alternately +-+-+-+-...+-+

OpenStudy (anonymous):

Ah good point. Well if you will notice the pattern that 1^3 - 99^3 is equivalent 2^3 + 98^3. You only have to divide your answer that you get from the equation in half.

OpenStudy (anonymous):

So instead of n^2(n+1)^2/4, you will have n^2(n+1)^2/8

OpenStudy (anonymous):

Sorry by equivalent i mean they canel each other out . . . wait a second is the answer 0? What do you think.

OpenStudy (anonymous):

in key answer book = 492500 (but didnt show steps) well, i checked by using your formula : n^2(n+1)^2/8 99^2 (100)^2 / 8 = 98010000/8 = 12251250 where the mistake ?

OpenStudy (anonymous):

I need help. Anyone out there see the problem?

OpenStudy (turingtest):

nobody seems to be acknowledging the alternating minus sign

OpenStudy (turingtest):

\[\sum_{n=1}^{99}(-1)^{n-1}n^3\]

OpenStudy (anonymous):

what's next, @TuringTest ? that still didnt make sense to me, but thanks trying helps me

OpenStudy (turingtest):

Well first, do you understand why I rewrote that series this way? do you see how the factor of (-1)^(n-1) will get the alternating +/- you have in your series?

OpenStudy (raden):

just alternative : it can rewitten as : A = 1^3 - 2^3 + 3^3 - 4^3 + 5^3 - 6^3 + ... + 99^3 A = 1^3+2^3+3^3+4^3+5^3+6^3+ ...+98^3+99^3 -2(2^3+4^3+6^3+...+98^3) A = 1^3+2^3+3^3+4^3+5^3+6^3+ ...+98^3+99^3 -2*2^3(1^3+2^3+3^3+...+49^3) A = 1^3+2^3+3^3+4^3+5^3+6^3+ ...+98^3+99^3 -16(1^3+2^3+3^3+...+49^3) A = (1+2+3+4+...+99)^2 - 16(1+2+3+4+...+49)^2 A = ..... continue it ...

OpenStudy (raden):

use the formula what @KirbyLegs said to calculating elements in paratheses 1+2+3+4+...+n = n(n+1)/2

OpenStudy (anonymous):

take the sum cube of odd numbers and subtract the sum of cubes of even number. That is, A = sum{1->50} (2k-1)^3 - sum{1->49} (2k)^3 All you got to do is to expand and use the formula of sum of k, k^2, k^3. The answer is 492,500. http://www.wolframalpha.com/input/?i=%28sum+k+%3D+1+to+50+%282k-1%29%5E3%29+-+%28sum+k+%3D+1+to+49+%282k%29%5E3%29

OpenStudy (anonymous):

A = (1+2+3+4+...+99)^2 - 16(1+2+3+4+...+49)^2 A = (99 * 100/2)^2 - 16(49 * 50/2)^2 A = 4950^2 - 16(1225)^2 A = 24502500 - 24010000 A = 492,500 waw, that's correct answer. Now make sense to me. Thank you very much every one :D

OpenStudy (raden):

you're welcome :*